OpenStudy (anonymous):

what is the quotient (x^3+8)/(x+2)

OpenStudy (anonymous):

its x^2-2x+4 (x-2)x+4 or x^2-2x+4

OpenStudy (yttrium):

as you can see. (x^3 + 8) = (x^3 + 2^3) = (x+2) (x^2 - 2x +4) Hence, (x^3+8)/(x+2) = [(x+2) (x^2 - 2x + 4)]/(x+2) So just solve for the final answer. :p

OpenStudy (anonymous):

Any of those are answers

OpenStudy (loser66):

@ohhh_itszoee

OpenStudy (anonymous):

Any of those are answers depending on how you look at it